Write the predecessor of: 
 (b) 10000
(c) 208090
(d) 7654321
 
eguwinwuraeriirhnuierghuyhiosdgjvnriuewghrn

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

The predecessors are


(b) 10000 – 1 = 9999

(c) 208090 – 1 = 208089

​(d) 7654321 – 1 = 7654320

Thanks